OpenStudy (anonymous):

Is x = 4 a solution to -3x – 12 = 0? Write yes or no.

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Substitute x=4 in the equation \[-3x-12=0\] \[-3\times 4-12=0\] \[-12-12=0\] \[-24 =0\] This is not true, so x=4 is not the solution of the equation

OpenStudy (anonymous):

I wouldn't say so.\[-3(4)-12=-24\neq0\]

OpenStudy (anonymous):

whoops nevermind